OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, 1) and (–2, 5)? A. 5/4 B. -5/4 C. -4/5 D. 4/5 I am pretty sure it is A right @pooja195 and @TheSmartOne

OpenStudy (anonymous):

Thanks!

OpenStudy (mathstudent55):

It's not A.

OpenStudy (mathstudent55):

The points are (3, 1) and (-2, 5). What are the y-coordinates?

TheSmartOne (thesmartone):

Wrong :P

TheSmartOne (thesmartone):

\(\sf\Large Slope = \frac{\color{blue}{y_2}-\color{red}{y_1}}{\color{green}{x_2}-\color{purple}{x_1}}\) Where the two points are in the form: \(\sf\Large (\color{purple}{x_1},\color{red}{y_1})~and~(\color{green}{x_2},\color{blue}{y_2})\)

OpenStudy (mathstudent55):

@Abbster2015 Let's do this correctly. Can you name the two y-coordinates of the given points?

OpenStudy (mathstudent55):

Here's a hint: The x-coordinates are in red, and the y-coordinates are in green. \(\large (\color{red}{x}, \color{green}{y})\) \(\large (\color{red}{3}, \color{green}{1})\) \(\large (\color{red}{-2}, \color{green}{5})\)

OpenStudy (anonymous):

I trust pooja because the slope formula is \[m = y2 - y1/x2 - x1\]

OpenStudy (mathstudent55):

Both The SmartOne and I have written that A is incorrect.

TheSmartOne (thesmartone):

correct, that is the formula but what are the x-ccoordinates and the y-coordinates? because your answer is wrong :P

OpenStudy (mathstudent55):

Pooja removed all her responses.

OpenStudy (anonymous):

Therefore the equation would be m= 3 + 2/5-1 so that would equal 5/4

OpenStudy (mathstudent55):

Let's do it together step by step and we'll find out what the correct answer is.

OpenStudy (anonymous):

Okay, I trust you guys.

TheSmartOne (thesmartone):

you're putting the x-coordinates on top, when they are at the bottom in the formula

OpenStudy (alexandervonhumboldt2):

the slope formula is \[\frac{ y_1-y_2 }{ x_1-x_2 }~~or~~\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (anonymous):

Okay.. so the x coordinates would be 5, 1?

OpenStudy (mathstudent55):

The y-coordinates are 1 and 5. In the same order, the x-coordinates are 3 and -2. We subtract the y-coordinates and the x-coordinates in the same order. \(\large m = \dfrac{5 - 1}{-2 - 3} \) @Abbster2015 Do you follwo so far?

TheSmartOne (thesmartone):

\(\sf\Large Slope = \frac{\color{blue}{y_2}-\color{red}{y_1}}{\color{green}{x_2}-\color{purple}{x_1}}\) Where the two points are in the form: \(\sf\Large (\color{purple}{3},\color{red}{1})~and~(\color{green}{-2},\color{blue}{5})\)

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (alexandervonhumboldt2):

5 and 1 are y coordinates

OpenStudy (mathstudent55):

The ***y-coordinates*** are 5 and 1.

OpenStudy (anonymous):

Yes I do, so that would equal 4/-5? Or -4/5

OpenStudy (mathstudent55):

The x-coordinates are -2 and 3.

TheSmartOne (thesmartone):

Now that is correct! @Abbster2015 :)

OpenStudy (mathstudent55):

\(\large m = \dfrac{5 - 1}{-2 - 3} = \dfrac{4}{-5} = -\dfrac{4}{5} \)

OpenStudy (mathstudent55):

As you can see, the answer is C, not A.

OpenStudy (anonymous):

So it would be C? Awesome thanks! WOuld you guys mind helping me with a few more? I don't mean to ask but this assissment is REALLY important. Like college type important.

TheSmartOne (thesmartone):

Sure :)

OpenStudy (anonymous):

Basiclly this is going to tell which college I go to *wipes sweat off forehead*

OpenStudy (anonymous):

Awesome I am going to open a new one.

OpenStudy (mathstudent55):

Remember that the slope is the difference in y divided by the difference in x. Also, remember that in an ordered pair, we list the x-coordinate before the y-coordinate. This is easy to remember, because an ordered pair is (x, y), where x and y are in alphabetical order.

OpenStudy (mathstudent55):

One last thing. You must find the difference in x and the difference in y using the coordinates in the same order.